When 9.0 g of hydrogen gas, H2, reacts with oxygen gas, O2, 73.0 g of water is produced.    2H2 + O2 --> 2H2O a) What is the theoretical yield of water? b) What is the percentage yield?


1
Expert's answer
2022-01-26T01:45:09-0500

Solution:

Calculate the moles of hydrogen gas (H2):

The molar mass of H2 is 2.0 g/mol

Hence,

Moles of H2 = (9.0 g H2) × (1 mol H2 / 2.0 g H2) = 4.5 mol H2


Balanced chemical equation:

2H2 + O2 → 2H2O

According to stoichiometry:

2 mol of H2 produces 2 mol of H2O

Thus, 4.5 mol of H2 produces:

(4.5 mol H2) × (2 mol H2O / 2 mol H2) = 4.5 mol H2O


Calculate the theoretical yield of water (H2O):

The molar mass of H2O is 18.0 g/mol

Hence,

Mass of H2O = (4.5 mol H2O) × (18.0 g H2O / 1 mol H2O) = 81.0 g H2O

The theoretical yield of water (H2O) is 81.0 grams


Calculate the percentage yield:

%yield = (actual yield / theoretical yield) × 100%

%yield = (73.0 g / 81.0 g) × 100% = 90.1%

The percentage yield is 90.12%


Answer:

The theoretical yield of water is 81.0 grams

The percentage yield is 90.12%